OpenStudy (anonymous):

I need your help with math, i need some examples of:: Commutative Property of Addition Associative Property of Addition Commutative Property of Multiplication Associative Property of Multiplication I totally forgot, and i guess my notes didn't save, or my mom deleted them!!

jimthompson5910 (jim_thompson5910):

Commutative Property of Addition a+b = b+a 2+3 = 3+2 1+4 = 4+1 ============================================= Associative Property of Addition a+(b+c) = (a+b)+c 3+(7+10) = (3+7)+10 9+(6+5) = (9+6)+5

jimthompson5910 (jim_thompson5910):

Commutative Property of Multiplication a*b = b*a 8*7 = 7*8 6*9 = 9*6 ============================================= Associative Property of Multiplication a*(b*c) = (a*b)*c 4*(5*8) = (4*5)*8 7*(2*9) = (7*2)*9

OpenStudy (anonymous):

Awesome!! thanks so much for helping me!!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

C: